# DualGR: Generative Retrieval with Long and Short-Term Interests Modeling

### v1

@@ Abstract.
In large-scale industrial recommendation systems, retrieval must produce high-quality candidates from massive corpora under strict latency. Recently, Generative Retrieval (GR) has emerged as a viable alternative to Embedding-Based Retrieval (EBR), which quantizes items into a finite token space and decodes candidates autoregressively, providing a scalable path that explicitly models target-history interactions via cross-attention. However, three challenges persist: 1) how to balance users’ long-term and short-term interests , 2) noise interference when generating hierarchical semantic IDs (SIDs), 3) the absence of explicit modeling for negative feedback such as exposed items without clicks. To address these challenges, we propose DualGR, a generative retrieval framework that explicitly models dual horizons of user interests with selective activation. Specifically, DualGR utilizes Dual-Branch Long/Short-Term Router (DBR) to cover both stable preferences and transient intents by explicitly modeling users’ long- and short-term behaviors. Meanwhile, Search-based SID Decoding (S2D) is presented to control context-induced noise and enhance computational efficiency by constraining candidate interactions to the current coarse (level-1) bucket during fine-grained (level-2/3) SID prediction. Finally, we propose an Exposure-aware Next-Token Prediction Loss (ENTP-Loss) that treats ”exposed-but-unclicked” items as hard negatives at level-1, enabling timely interest fade-out. On the large-scale Kuaishou short-video recommendation system, DualGR has achieved outstanding performance. Online A/B testing shows +0.527% video views and +0.432% watch time lifts, validating DualGR as a practical and effective paradigm for industrial generative retrieval.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Retrieval models and ranking
@@ 1. Introduction
In large-scale industrial recommender systems, retrieval is the first stage, which must select thousands of high-quality candidates from a multi-billion-item corpus for downstream ranking. Our work is deployed in Kuaishou’s double-column Explore feed, akin to Instagram’s two-column waterfall layout, where multiple cards appear on the same screen. Compared with the single-column full-screen feed, this layout places stricter demands on candidate diversity, which must produce diverse yet relevant candidates in real time to cover multiple topics and intents, also to rapidly track emerging interests, therefore it magnifies modeling and efficiency challenges at the retrieval stage. In industry, the dominant approach is Embedding-Based Retrieval (EBR), whose DNN twin-tower architecture retrieves user-item matches via approximate nearest neighbors (ANN)  (meng2025user; liu2024kuaiformer). While highly scalable, EBR is inherently limited in multi-interest modeling and target-aware interaction, and is vulnerable to exposure bias (cen2020controllable; yi2019sampling).
@@
Recently, inspired by the generative progress of LLMs and Transformers (zhao2023survey), industry has begun adopting Generative Retrieval (GR), whose items are discretized into multi-level semantic IDs (SIDs) and candidates are decoded autoregressively along the hierarchy (rajput2023recommender; yang2024unifying; zhou2025onerec). Operating within the same finite and discrete token space during both training and inference can effectively alleviate exposure bias; moreover, cross-attention in decoding enables explicit target-history interaction. Combined with beam search, the model can naturally produce diverse candidate sets under comparable latency budgets, aligning with retrieval’s coverage and exploration goals. Despite this promise, deploying the generative paradigm robustly in multi-interest, short-video retrieval still faces three key challenges: ❶ Long-short interest interference. The double-column feed is highly sensitive to both persistent preferences and hotspot tracking; existing ”multi-interest” effects often emerge passively via multi-beam decoding and lack explicit control across time scales. When stable preferences and transient hotspots coexist, attention and gradients can dilute each other, destabilizing routing and decoding and upsetting the relevance-diversity balance. ❷ Context noise and long-history constraints. During fine-grained (level-2/3) SIDs generation, the model struggles to separate target-relevant fragments from irrelevant noise, akin to context engineering issues observed in LLMs (liu2023lost). In the same time, under the strict latency budgets of retrieval, without category-wise coarse screening before interaction, the use of long histories will become prohibitive (pi2020search; meng2025user). ❸ Missing negative feedback. Industrial logs contain many unclicked exposed items, yet in the next-token prediction (NTP) (rajput2023recommender; zhou2025onerec), it’s difficult to encode them as explicit negative signals, impeding the timely fade-out of non-interest directions and thereby degrading decoding quality and coverage efficiency.
@@
To address these issues, we propose DualGR, a generative retrieval framework with two explicit interest branches and search-based decoding, while enabling the timely fade-out of non-interest. First, we utilizes Dual-Branch Long/Short-Term Router (DBR) to decompose a user’s history into long-term and short-term summaries, compute target-aware similarities, and route to exactly one branch to suppress cross-branch dilution; at inference, run one forward pass per branch and merge the candidates to cover both stable preferences and transient intents. Second, we introduce Search-based SID Decoding (S2D), when predicting fine-grained (level-2/3) SIDs, candidate actions are constrained to the current coarse (level-1) bucket, which shrinks the search space and reinforces intra-bucket consistency, thereby enabling longer usable histories while reducing noise interference and improving computational efficiency. Third, to overcome the lack of negatives, we propose Exposure-aware Next-Token Prediction Loss (ENTP-Loss), at level-1, we treat unclicked exposures as hard negatives, explicitly penalizing the corresponding coarse categories to accelerate the fade-out of non-interest directions. Considering latency and compute budgets, we adopt an encoder-free lightweight frontend (embeddings + LayerNorm) and perform step-wise cross attention over the action list, achieving SIM-like target awareness without redundant re-encoding (pi2020search). The contributions are as follows:
@@ 2.1. Preliminary
For generative retrieval, each video $v\in\mathcal{V}$ is discretized into hierarchical semantic IDs (SIDs): $\big(s^{(1)},\,s^{(2)},\,\ldots,\,s^{(L)}\big)$, where $s^{(\ell)}\in\mathcal{C}^{(\ell)}$, $\mathcal{C}^{(\ell)}$ is the level-$\ell$ codebook  (rajput2023recommender; yang2024unifying; zhou2025onerec). A generative retriever defines a conditional distribution over hierarchical SIDs via:
@@
At serving time, the Top-$K$ SIDs are decoded under beam search (beam size $B$) or sampling (zhou2025onerec): $\widehat{\mathcal{S}}_{\theta}(x_{t})=\operatorname{TopK}_{B}\Big\{p_{\theta}\!\big(s^{(1:L)}\!\mid x_{t}\big)\Big\}.$ Finally, a prebuilt mapping/index $\mathrm{Map}:\mathcal{S}\rightarrow\mathcal{V}$ projects SIDs back to concrete video ids forming the retrieval result $\mathcal{R}_{\theta}(x_{t})=\mathrm{Map}\!\big(\widehat{\mathcal{S}}_{\theta}(x_{t})\big)\subseteq\mathcal{V}$, which is also required to satisfy Eq. 1.
@@ 2.2. Framework Overview
We first pre-quantize videos into L=3L{=}3 hierarchical SIDs via residual k-means (RQ-KMeans) (zhou2025onerec). Built on an encoder-free, decoder-centric backbone, DualGR integrates three components, i.e., DBR (long/short routing), S2D (search-based fine decoding), and ENTP-Loss (exposure-aware training) as illustrated in Fig. 1.
@@ 2.4.1. Training
then builds the fine-level input by concatenating user static features with the selected actions and encoding them as in Eq. 6 to obtain $x_{t}^{(>1)}$. Under teacher forcing (zhao2023survey), the conditional likelihood for the remaining tokens is
@@ 2.4.2. Inference
S2D realizes the decoding with searching, controlled noise and latency. Conditioning on the chosen level-1 bucket $s_{\star}^{(1)}$, the size of the in-bucket history scales approximately as $\big|\mathcal{H}_{t}\!\big(s_{\star}^{(1)}\big)\big|\approx\frac{|\mathcal{H}_{t}|}{|\mathcal{V}^{(1)}|}$. In practice, the codebook per level is typically large (e.g., $\geq 256$ and up to 81928192(rajput2023recommender; zhou2025onerec), hence $\big|\mathcal{H}_{t}(s_{\star}^{(1)})\big|\!\ll\!|\mathcal{H}_{t}|$. This shrinkage enables longer usable histories under a fixed latency budget.
@@ 3. Experiments
Experimental Settings. We conduct experiments on Kuaishou’s double-column Explore feed short-video stream, one of the platform’s largest recommendation surfaces, serving over 100M users, 1B video views, and 10B exposures per day. We compare DualGR against strong retrieval baselines deployed in this scenario, including ComiRec (cen2020controllable), PDN (li2021path), and KuaiFormer (liu2024kuaiformer); additionally, for a fair comparison with generative retrieval, we implement TIGER (rajput2023recommender). For evaluation, we use Online Hit Rate (HR), a metric widely adopted in prior work (liu2024kuaiformer). DualGR adopts an encoder-free decoder of 4 Transformer blocks with model dimension 512, 8 attention heads, and feed-forward dimension 1024. Videos are pre-quantized into L=3L{=}3 hierarchical codebooks with per-level size $|\mathcal{C}^{(\ell)}|{=}8192$. For inputs, we use the user’s most recent positive interactions where $|\mathcal{H}_{t}|{=}1000$, the DBR employs window lengths $L_{\text{long}}{=}1000$ and $L_{\text{short}}{=}64$. Training uses the ENTP-Loss with $\alpha{=}0.1$.

### v2

@@ Abstract.
In large-scale industrial recommendation systems, retrieval must produce high-quality candidates from massive corpora under strict latency. Recently, Generative Retrieval (GR) has emerged as a viable alternative to Embedding-Based Retrieval (EBR), which quantizes items into a finite token space and decodes candidates autoregressively, providing a scalable path that explicitly models target-history interactions via cross-attention. However, deploying GR in short-video feeds remains challenged by long-short interest interference, context-induced noise in hierarchical SID generation, and the lack of explicit learning from exposed-but-unclicked feedback. To address these challenges, we propose DualGR, which combines (i) a Dual-Branch Long/Short-Term Router (DBR) with selective activation, (ii) Search-based SID Decoding (S2D) that constrains fine-level decoding within the current coarse bucket for efficiency and noise control, and (iii) an Exposure-aware Next-Token Prediction Loss (ENTP-Loss) that treats unclicked exposures as coarse-level hard negatives to promote timely interest fade-out. On the large-scale Kuaishou short-video recommendation system, DualGR has achieved outstanding performance. Online A/B testing shows +0.527% video views and +0.432% watch time lifts, validating DualGR as a practical and effective paradigm for industrial generative retrieval.
@@
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emiratesbooktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792851isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Retrieval models and ranking
@@ 1. Introduction
In large-scale industrial recommender systems, retrieval is the first stage, which must select thousands of high-quality candidates from a multi-billion-item corpus for downstream ranking. Our work is deployed in Kuaishou’s double-column Explore feed, akin to Instagram’s two-column waterfall layout, where multiple cards appear on the same screen. Compared with the single-column full-screen feed, this layout places stricter demands on candidate diversity, which must produce diverse yet relevant candidates in real time to cover multiple topics and intents, also to rapidly track emerging interests, therefore it magnifies modeling and efficiency challenges at the retrieval stage. In industry, the dominant approach is Embedding-Based Retrieval (EBR), whose DNN twin-tower architecture retrieves user-item matches via approximate nearest neighbors (ANN)  (Meng et al., 2025; Liu et al., 2024). While highly scalable, EBR is inherently limited in multi-interest modeling and target-aware interaction, and is vulnerable to exposure bias (Cen et al., 2020; Yi et al., 2019).
@@
Recently, inspired by the generative progress of LLMs and Transformers (Zhao et al., 2023), industry has begun adopting Generative Retrieval (GR), whose items are discretized into multi-level semantic IDs (SIDs) and candidates are decoded autoregressively along the hierarchy (Rajput et al., 2023; Yang et al., 2024; Zhou et al., 2025). Operating within the same finite and discrete token space during both training and inference can effectively alleviate exposure bias; moreover, cross-attention in decoding enables explicit target-history interaction. Combined with beam search, the model can naturally produce diverse candidate sets under comparable latency budgets, aligning with retrieval’s coverage and exploration goals. Despite this promise, deploying the generative paradigm robustly in multi-interest, short-video retrieval still faces three key challenges: ❶ Long-short interest interference. The double-column feed is highly sensitive to both persistent preferences and hotspot tracking; existing ”multi-interest” effects often emerge passively via multi-beam decoding and lack explicit control across time scales. When stable preferences and transient hotspots coexist, attention and gradients can dilute each other, destabilizing routing and decoding and upsetting the relevance-diversity balance. ❷ Context noise and long-history constraints. During fine-grained (level-2/3) SIDs generation, the model struggles to separate target-relevant fragments from irrelevant noise, akin to context engineering issues observed in LLMs (Liu et al., 2023). In the same time, under the strict latency budgets of retrieval, without category-wise coarse screening before interaction, the use of long histories will become prohibitive (Pi et al., 2020; Meng et al., 2025). ❸ Missing negative feedback. Industrial logs contain many unclicked exposed items, yet in the next-token prediction (NTP) (Rajput et al., 2023; Zhou et al., 2025), it’s difficult to encode them as explicit negative signals, impeding the timely fade-out of non-interest directions and thereby degrading decoding quality and coverage efficiency.
@@
To address these issues, we propose DualGR, a generative retrieval framework with two explicit interest branches and search-based decoding, while enabling the timely fade-out of non-interest. First, we utilizes Dual-Branch Long/Short-Term Router (DBR) to decompose a user’s history into long-term and short-term summaries, compute target-aware similarities, and route to exactly one branch to suppress cross-branch dilution; at inference, run one forward pass per branch and merge the candidates to cover both stable preferences and transient intents. Second, we introduce Search-based SID Decoding (S2D), when predicting fine-grained (level-2/3) SIDs, candidate actions are constrained to the current coarse (level-1) bucket, which shrinks the search space and reinforces intra-bucket consistency, thereby enabling longer usable histories while reducing noise interference and improving computational efficiency. Third, to overcome the lack of negatives, we propose Exposure-aware Next-Token Prediction Loss (ENTP-Loss), at level-1, we treat unclicked exposures as hard negatives, explicitly penalizing the corresponding coarse categories to accelerate the fade-out of non-interest directions. Considering latency and compute budgets, we adopt an encoder-free lightweight frontend (embeddings + LayerNorm) and perform step-wise cross attention over the action list, achieving SIM-like target awareness without redundant re-encoding (Pi et al., 2020). The contributions are as follows:
@@ 2.1. Preliminary
For generative retrieval, each video $v\in\mathcal{V}$ is discretized into hierarchical semantic IDs (SIDs): $\big(s^{(1)},\,s^{(2)},\,\ldots,\,s^{(L)}\big)$, where $s^{(\ell)}\in\mathcal{C}^{(\ell)}$, $\mathcal{C}^{(\ell)}$ is the level-$\ell$ codebook  (Rajput et al., 2023; Yang et al., 2024; Zhou et al., 2025). A generative retriever defines a conditional distribution over hierarchical SIDs via:
@@
At serving time, the Top-$K$ SIDs are decoded under beam search (beam size $B$) or sampling (Zhou et al., 2025): $\widehat{\mathcal{S}}_{\theta}(x_{t})=\operatorname{TopK}_{B}\Big\{p_{\theta}\!\big(s^{(1:L)}\!\mid x_{t}\big)\Big\}.$ Finally, a prebuilt mapping/index $\mathrm{Map}:\mathcal{S}\rightarrow\mathcal{V}$ projects SIDs back to concrete video ids forming the retrieval result $\mathcal{R}_{\theta}(x_{t})=\mathrm{Map}\!\big(\widehat{\mathcal{S}}_{\theta}(x_{t})\big)\subseteq\mathcal{V}$, which is also required to satisfy Eq. 1.
@@ 2.2. Framework Overview
We first pre-quantize videos into L=3L{=}3 hierarchical SIDs via residual k-means (RQ-KMeans) (Zhou et al., 2025). Built on an encoder-free, decoder-centric backbone, DualGR integrates three components, i.e., DBR (long/short routing), S2D (search-based fine decoding), and ENTP-Loss (exposure-aware training) as illustrated in Fig. 1.
@@ 2.4.1. Training
then builds the fine-level input by concatenating user static features with the selected actions and encoding them as in Eq. 6 to obtain $x_{t}^{(>1)}$. Under teacher forcing (Zhao et al., 2023), the conditional likelihood for the remaining tokens is
@@ 2.4.2. Inference
S2D realizes the decoding with searching, controlled noise and latency. Conditioning on the chosen level-1 bucket $s_{\star}^{(1)}$, the size of the in-bucket history scales approximately as $\big|\mathcal{H}_{t}\!\big(s_{\star}^{(1)}\big)\big|\approx\frac{|\mathcal{H}_{t}|}{|\mathcal{V}^{(1)}|}$. In practice, the codebook per level is typically large (e.g., $\geq 256$ and up to 81928192(Rajput et al., 2023; Zhou et al., 2025), hence $\big|\mathcal{H}_{t}(s_{\star}^{(1)})\big|\!\ll\!|\mathcal{H}_{t}|$. This shrinkage enables longer usable histories under a fixed latency budget.
@@ 3. Experiments
Experimental Settings. We conduct experiments on Kuaishou’s double-column Explore feed short-video stream, one of the platform’s largest recommendation surfaces, serving over 100M users, 1B video views, and 10B exposures per day. We compare DualGR against strong retrieval baselines deployed in this scenario, including ComiRec (Cen et al., 2020), PDN (Li et al., 2021), and KuaiFormer (Liu et al., 2024); additionally, for a fair comparison with generative retrieval, we implement TIGER (Rajput et al., 2023). For evaluation, we use Online Hit Rate (HR), a metric widely adopted in prior work (Liu et al., 2024). DualGR adopts an encoder-free decoder of 4 Transformer blocks with model dimension 512, 8 attention heads, and feed-forward dimension 1024. Videos are pre-quantized into L=3L{=}3 hierarchical codebooks with per-level size $|\mathcal{C}^{(\ell)}|{=}8192$. For inputs, we use the user’s most recent positive interactions where $|\mathcal{H}_{t}|{=}1000$, the DBR employs window lengths $L_{\text{long}}{=}1000$ and $L_{\text{short}}{=}64$. Training uses the ENTP-Loss with $\alpha{=}0.1$.
@@ 4. Conclusion
Acknowledgements
This paper is partially supported by the National Natural Science Foundation of China (No.62502488, No.12227901), Natural Science Foundation of Jiangsu Province (BK20240460), the grant from State Key Laboratory of Resources and Environmental Information System.